Addition, Subtraction, & Complex Fractions
(6x + 1)/(x^2 - 9) + (4)/(x - 3)
my best attempt is:
([6x + 1]*[x - 3]) * ([x^2 - 9]*[4]) / ([x^2 - 9]*[x - 3])

Answer by richwmiller(17219)

x^2-9 = (x-3)*(x+3)
so you get
((6x+1)+4*(x+3))/(x^2-9)
6x+1+4x+12/(x^2-9)=10x+13/(x^2-9)
(10x+13)/(x-3)*(x+3)